if a+b+c=12 and a^2+b^2+c^2=11 find ab+bc+ca
And (A+B+C) a^2+b^2+c^2-ab-bc-ca
 
 

Grade:9

1 Answers

Vikas TU
4 years ago
Dear student 
Dear student 
(a+b+c)^2 = a^2+b^2+c^2+2ab+2bc+2ca (formula)
(a+b+c)^2 = a^2 + b^2+ c^2 + 2(ab+bc+ca)
144 = 11 + 2(ab+bc+ca)
 (A+B+C) a^2+b^2+c^2-ab-bc-ca = 12(11) – 133/2 
Hope this helps 
Good Luck 

133/2 = (ab+bc+ca)